# Stepwise Guided Policy Optimization: Coloring your Incorrect Reasoning in GRPO

### v4

@@ Appendix A Related Works
Reinforcement learning from human feedback (RLHF) uses human-preference-aligned reward models to evaluate response quality [Christiano et al., 2017, Ziegler et al., 2019, Stiennon et al., 2020, Ouyang et al., 2022]. A key barrier to scale RLHF is the need for high-quality human labels. Previous studies [Gilardi et al., 2023, Ding et al., 2023] have shown that modern LLMs exhibit strong alignment with human judgments, suggesting that AI-generated labels can serve as a viable alternative. In this context, [Bai et al., 2022] was the first to explore RLAIF, jointly optimizing helpfulness and harmlessness using both human and AI-generated labels, and [Roit et al., 2023, Kwon et al., 2023, Lee et al., 2024] showed that LLMs can produce informative reward signals for RL post-training. Our approach can leverage AI feedback to introduce response diversity within all-negative-sample groups by assigning intermediate binary rewards to reasoning steps. Indeed, one identifies the proportion of correct steps in the reasoning trajectory and compute a reward $r_{i}\in[0,1)$.
Appendix B Additional Experimental Results
Beyond aggregate results, we provide a targeted analysis of SGPO’s impact. In line with our theoretical finding that SGPO converges faster than GRPO, empirical metrics offer supporting evidence. Prior work on RLVR entropy highlights its link to performance: Cui et al. [2025] showed that lower policy entropy under correct signals correlates with stronger policies, while Agarwal et al. [2025] demonstrated that directly minimizing entropy can improve performance. As shown in Figure 3, SGPO reduces policy entropy more rapidly than GRPO, indicating faster convergence toward deterministic RLVR behavior with higher rollout confidence. This matches our theoretical results, confirming that step-wise signals accelerate convergence.
Appendix C Missing Proofs
We first present the detailed setup for our stylized model and prove several technical lemmas. Then, we use these lemmas to prove our main result in Theorem 3.3.
C.1 Stylized model
We consider a policy parameterized by a softmax function, which is standard in the analysis of reinforcement learning methods [Agarwal et al., 2020, Mei et al., 2021, Li et al., 2024b]:
$\pi_{\theta}(a_{1:T}\,|\,\mathbf{x})=\prod_{t=1}^{T}\pi_{\theta_{t}}(a_{t}\,|\,\mathbf{x},a_{1:t-1})=\prod_{t=1}^{T}\tfrac{\exp(\theta_{t}^{\mathbf{x},a_{1:t-1},a_{t}})}{\sum_{a^{\prime}_{t}\in\mathcal{V}^{\star}}\exp(\theta_{t}^{\mathbf{x},a_{1:t-1},a_{t}^{\prime}})},$
By convention, we assume that $\pi_{\theta_{1}}(a_{1}\,|\,\mathbf{x},a_{1:0})=\pi_{\theta_{1}}\left(a_{1}\,|\,\mathbf{x}\right)$.
For simplicity, we perform our analysis in the likelihood space rather than in the parameter space (i.e., $\theta$) directly. Indeed, we define the key quantities as follows,
$p\doteq\pi_{\theta_{1}}(a_{1}=2\,|\,\mathbf{x})=\tfrac{e^{\theta_{1}^{\mathbf{x},2}}}{e^{\theta_{1}^{\mathbf{x},1}}+e^{\theta_{1}^{\mathbf{x},2}}},\qquad q\doteq\pi_{\theta_{2}}(a_{2}=2\,|\,\mathbf{x},a_{1}=2)=\tfrac{e^{\theta_{2}^{\mathbf{x},2,2}}}{e^{\theta_{2}^{\mathbf{x},2,1}}+e^{\theta_{2}^{\mathbf{x},2,2}}}.$
We rewrite the generic GRPO update with a step size $\eta>0$ as follows,
where $N$ is the number of prompts, $G$ is the number of groups, $H$ is the number of reasoning steps in each response, $s_{\theta}(\mathbf{x}^{i},a^{i,k}_{1:h-1}):=\nabla_{\theta}\log\pi_{\theta}(a_{t}|\mathbf{x},a_{1:h-1})$ is the score function, and the advantage $A_{i,k}$ is defined by
$A_{i,k}=\tfrac{r(\mathbf{x}^{i},\mathbf{y}^{i,k})-(1/G)\sum_{j=1}^{G}r(\mathbf{x}^{i},\mathbf{y}^{i,j})}{\sqrt{(1/G)\sum_{j=1}^{G}(r(\mathbf{x}^{i},\mathbf{y}^{i,j})-(1/G)\sum_{j^{\prime}=1}^{G}r(\mathbf{x}^{i},\mathbf{y}^{i,j^{\prime}}))^{2}}}$
To distinguish, we denote $g_{\texttt{GRPO}}(\cdot)$ as the gradient estimator using classical outcome reward model $r$, and $g_{\texttt{SGPO}}(\cdot)$ as the gradient estimator using the reward $r_{\texttt{SGPO}}$ as proposed in Section 3.1.
and
Note that we restrict the sample space to $\mathbf{y}\in\{(1,1),(2,1),(2,2)\}$, excluding the sample $(1,2)$. The responses can be drawn i.i.d. from the distribution as follows,
We set $G=2$ and focus on the SGPO and GRPO training dynamics with population-level policy gradient which can be computed exactly for the stylized model as follows,
Since $g_{\texttt{GRPO}}(\theta)$ and $g_{\texttt{SGPO}}(\theta)$ concentrate around $\bar{g}_{\texttt{GRPO}}(\theta)$ and $\bar{g}_{\texttt{SGPO}}(\theta)$ when the number of samples in each group is sufficiently large, it is reasonable to analyze the population-level dynamics at first. Note that the high-probability guarantees for the sample-level dynamics can be derived using concentration inequalities under certain conditions.
We can explicitly write down the SGPO and GRPO update rules with $\eta=1$ using the likelihood parameters $p$ and $q$ as follows,
where the functions $f_{ij}$ are defined by
$f_{11}(p)=\log(p)+p(1-p)-\log(1-p+pe^{p(1-p)}),$ (4)
$f_{21}(p,q)=\log(p)+p(1-p)q-\log(1-p+pe^{p(1-p)q}),$
$f_{12}(p,q)=\log(q)+p^{2}q(1-q)-\log(1-q+qe^{p^{2}q(1-q)}),$
$f_{22}(p,q)=\log(q)+pq(1-q)-\log(1-q+qe^{pq(1-q)}).$
C.2 Technical lemmas
@@
We provide several technical lemmas that are important to the subsequent proof of Theorem 3.3. Indeed, the first lemma summarizes the properties of particular functions related to the aforementioned functions $f_{11}$, $f_{21}$, $f_{12}$ and $f_{22}$ from Eq. 4.
Lemma C.1.
@@
    The function $\varphi(x):=\log(1+(1/2)e^{-e^{x}})$ is strictly concave on $(-\infty,0)$.
@@
$f^{\prime}_{11}(p_{1})=\tfrac{1+(1-2p_{1})p_{1}(1-p_{1})}{p_{1}(1+p_{1}+p_{1}e^{p_{1}(1-p_{1})})}>\tfrac{3}{4p_{1}(1+p_{1}+p_{1}e^{p_{1}(1-p_{1})})}>0.$
@@
$\tfrac{\partial f_{21}(p_{1},p_{2})}{p_{1}}$ $=$ $\tfrac{1+p_{2}(1-2p_{1})p_{1}(1-p_{1})}{p_{1}(1+p_{1}+p_{1}e^{p_{2}p_{1}(1-p_{1})})}\ >\ \tfrac{3}{4p_{1}(1+p_{1}+p_{1}e^{p_{1}(1-p_{1})})}>0,$
$\tfrac{\partial f_{21}(p_{1},p_{2})}{p_{2}}$ $=$ $\tfrac{p_{1}(1-p_{1})^{2}}{1+p_{1}+p_{1}e^{p_{2}p_{1}(1-p_{1})}}\ >\ 0.$
@@
$\varphi^{\prime\prime}(x)=\tfrac{(e^{x+e^{x}}/2-e^{e^{x}}/2-1/4)e^{x}}{e^{2e^{x}}+e^{e^{x}}+1/4}.$
@@
$(ue^{u}/2-e^{u}/2-1/4)u=(e^{u}(u-1)/2-1/4)u<-(1/4)u<0.$
Thus, $\varphi^{\prime\prime}(x)<0$ for all $x<0$ which shows that $f$ is strictly concave on $(-\infty,0)$. $\Box$
@@
By Lemma C.3, this is indeed true. This completes the proof. $\Box$
Lemma C.3.
@@
where $c>0$ is a universal constant, and single-variable polynomials $c_{20}(s),\ldots,c_{2}(s),c_{0}(s)>0$ and $\Delta_{2}:=c_{1}(s)^{2}-4c_{2}(s)c_{0}(s)<0$, for all $s>0$ (see Section C.2 for details). Notice that from the table, we can see the only nontrivial parts are $c_{3}(s)>0$ and $c_{2}(s)>0$ because only these two contain negative coefficients. The positivity of $c_{3}(s)$ is simple because there is only one term ($s^{9}$) with negative coefficient and for all $s>0$,
@@
Finally, we have $p_{\texttt{SGPO}}^{(0)}\geq q_{\texttt{SGPO}}^{(0)}$. Thus, it suffices to show that $p_{\texttt{SGPO}}^{(k)}\geq q_{\texttt{SGPO}}^{(k)}$ implies $p_{\texttt{SGPO}}^{(k+1)}>q_{\texttt{SGPO}}^{(k+1)}$ for all $k\geq 0$. Indeed, Lemma C.1(i) and $p_{\texttt{SGPO}}^{(k)}\geq q_{\texttt{SGPO}}^{(k)}$ yield
@@
Then, Lemma C.1(ii) and $p_{\texttt{SGPO}}^{(k)}\in(0,1)$ yield
@@
C.3 Proof of Theorem 3.3
To show (i), recall that the sequence $(p_{\texttt{SGPO}}^{(k)})_{k\in\mathbb{N}}$ is strictly increasing and bounded in $(0,1)$ from Lemmas C.4(i) and C.4(ii), so it converge to some value $c\in(0,1]$. Take limit as $k\to\infty$:
$1=\lim_{k\to\infty}\tfrac{p_{\texttt{SGPO}}^{(k+1)}}{p_{\texttt{SGPO}}^{(k)}}=\lim_{k\to\infty}\tfrac{1}{(1-p_{\texttt{SGPO}}^{(k)})e^{-\Delta_{\texttt{SGPO},p}^{(k)}}+p_{\texttt{SGPO}}^{(k)}}=\tfrac{1}{(1-c)e^{-c(1-c)}+c}.$
Using the simple Taylor lower bound $e^{-x}\geq 1-x$, we have
To show (ii), consider the base case:
where the inequality follows from Lemma C.1(ii). Thus, we use induction and assume $p_{\texttt{SGPO}}^{(k)}>p_{\texttt{GRPO}}^{(k)}$ for some $k\geq 1$. Then we have,
where the first inequality uses Lemma C.1(i) and the second one uses Lemma C.1(ii). Thus, $p_{\texttt{SGPO}}^{(k+1)}>p_{\texttt{GRPO}}^{(k+1)}$ and induction completes. We have proved that $p_{\texttt{SGPO}}^{(k)}>p_{\texttt{GRPO}}^{(k)}$ for all $k\geq 1$.
Also recall the update rule of $p_{\texttt{SGPO}}^{(k)}$ and $q_{\texttt{SGPO}}^{(k)}$:
$\begin{array}[]{lcl}p_{\texttt{SGPO}}^{(k+1)}&=&\tfrac{1}{(1/p_{\texttt{SGPO}}^{(k)}-1)\exp(-p_{\texttt{SGPO}}^{(k)}(1-p_{\texttt{SGPO}}^{(k)}))+1}\\ q_{\texttt{SGPO}}^{(k+1)}&=&\tfrac{1}{(1/q_{\texttt{SGPO}}^{(k)}-1)\exp(-(p_{\texttt{SGPO}}^{(k)})^{2}q_{\texttt{SGPO}}^{(k)}(1-q_{\texttt{SGPO}}^{(k)}))+1}\end{array},$
and it suffices to show $p_{\texttt{SGPO}}^{(k)}q_{\texttt{SGPO}}^{(k)}>(p_{\texttt{GRPO}}^{(k)})^{2}$ for all $k\geq 1$. We prove by induction. For the base case,
$\sqrt{p_{\texttt{SGPO}}^{(1)}q_{\texttt{SGPO}}^{(1)}}=\sqrt{\tfrac{1}{1+(1/2)e^{-1/4}}\cdot\tfrac{1}{1+(1/2)e^{-1/16}}}>\tfrac{1}{1+(1/2)e^{-1/8}}=p_{\texttt{GRPO}}^{(1)}.$
The above inequality holds true since Lemma C.1(iv) implies
$2\log(1+(1/2)e^{-1/8})>\log(1+(1/2)e^{-1/4})+\log(1+(1/2)e^{-1/16}),$
It remains to show that $p_{\texttt{SGPO}}^{(k)}q_{\texttt{SGPO}}^{(k)}>(p_{\texttt{GRPO}}^{(k)})^{2}$ implies $p_{\texttt{SGPO}}^{(k+1)}q_{\texttt{SGPO}}^{(k+1)}>(p_{\texttt{GRPO}}^{(k+1)})^{2}$ for $k\geq 1$. By Lemma C.4(iii), we know $p_{\texttt{SGPO}}^{(k)}>q_{\texttt{SGPO}}^{(k)}$ for all $k\geq 1$. Thus, Lemma C.2 implies that
Using Lemma C.1(iii), we complete the induction by applying our induction hypothesis:
$\tfrac{1}{C\left(\sqrt{p_{\texttt{SGPO}}^{(k)}q_{\texttt{SGPO}}^{(k)}}\right)^{2}}=\exp\left(f_{21}\left(\sqrt{p_{\texttt{SGPO}}^{(k)}q_{\texttt{SGPO}}^{(k)}},\sqrt{p_{\texttt{SGPO}}^{(k)}q_{\texttt{SGPO}}^{(k)}}\right)\right)>\exp(f_{21}(p_{\texttt{GRPO}}^{(k)},p_{\texttt{GRPO}}^{(k)}))=(p_{\texttt{GRPO}}^{(k+1)})^{2}.$
This completes the proof.
C.4 Empirical evaluations
We conduct simple simulations to compare SGPO and GRPO under the stylized model and plot the resulting learning curves in Figure 4. The left panel shows the probability of selecting the “good” action in the first step at iteration $k$ (i.e., $p_{\texttt{SGPO}}^{(k)}$ vs. $p_{\texttt{GRPO}}^{(k)}$), while the right panel shows the probability of learning the optimal policy (i.e., $p_{\texttt{SGPO}}^{(k)}q_{\texttt{SGPO}}^{(k)}$ vs. $p_{\texttt{GRPO}}^{(k)}q_{\texttt{GRPO}}^{(k)}$). The results align with the predictions of Theorem 3.3, demonstrating that the likelihood of learning the optimal policy under SGPO consistently exceeds that of GRPO across training.

### v5

@@ Appendix A Related Works
Reinforcement learning from human feedback (RLHF) uses human-preference-aligned reward models to evaluate response quality [Christiano et al., 2017, Ziegler et al., 2019, Stiennon et al., 2020, Ouyang et al., 2022]. A key barrier to scale RLHF is the need for high-quality human labels. Previous studies [Gilardi et al., 2023, Ding et al., 2023, Zhang et al., 2024b] have shown that modern LLMs exhibit strong alignment with human judgments, suggesting that AI-generated labels can serve as a viable alternative. In this context, [Bai et al., 2022] was the first to explore RLAIF, jointly optimizing helpfulness and harmlessness using both human and AI-generated labels, and [Roit et al., 2023, Kwon et al., 2023, Lee et al., 2024] showed that LLMs can produce informative reward signals for RL post-training. Our approach can leverage AI feedback to introduce response diversity within all-negative-sample groups by assigning intermediate binary rewards to reasoning steps. Indeed, one identifies the proportion of correct steps in the reasoning trajectory and use it to compute a reward $r_{i}\in[0,1)$.
Appendix B Missing Proofs
In this section, we present the derivations underlying the update rules in Section 3.2. In particular, we derive $g_{\texttt{GRPO}}(\theta)$ and $g_{\texttt{SGPO}}(\theta)$, and obtain the corresponding update forms in Eqs. 3 and 4.
To derive $g_{\texttt{GRPO}}(\theta)$ and $g_{\texttt{SGPO}}(\theta)$, we start by restating the score functions:
$s(a_{1}=1\mid x)=\begin{bmatrix}p\\ -p\\ 0\\ 0\end{bmatrix},\quad s(a_{1}=2\mid x)=\begin{bmatrix}p-1\\ 1-p\\ 0\\ 0\end{bmatrix}$,
$s(a_{2}=1\mid x,a_{1}=2)=\begin{bmatrix}0\\ 0\\ q\\ -q\end{bmatrix},\quad s(a_{2}=2\mid x,a_{1}=2)=\begin{bmatrix}0\\ 0\\ q-1\\ 1-q\end{bmatrix}$.
Let $S(y):=\sum_{h=1}^{H}s(a_{h}\mid x,a_{1:h-1})$ denote the score-sum for trajectory $y=(a_{1},a_{2})$. Under the restricted trajectory space $\{(1,1),(2,1),(2,2)\}$, we have
$S(1,1)=\begin{bmatrix}p\\ -p\\ 0\\ 0\end{bmatrix},\quad S(2,1)=\begin{bmatrix}p-1\\ 1-p\\ q\\ -q\end{bmatrix},\quad S(2,2)=\begin{bmatrix}p-1\\ 1-p\\ q-1\\ 1-q\end{bmatrix}$,
and the trajectory probabilities are
$\mathbb{P}(1,1)=1-p,\quad\mathbb{P}(2,1)=p(1-q),\quad\mathbb{P}(2,2)=pq.$
Recall $G=2$ and $H=2$. For a single prompt, our estimator averages over both samples and steps:
$g(\theta)=\tfrac{1}{GH}\left(\sum_{k=1}^{G}\sum_{h=1}^{H}s\left(a^{(k)}_{h}\mid x,a^{(k)}_{1:h-1}\right)A_{k}\right)=\tfrac{1}{GH}\left(\sum_{k=1}^{G}S(y^{(k)})A_{k}\right),$
where $A_{k}$ is the within-group standardized advantage computed from the two rewards. For $G=2$, whenever the two rewards are distinct, standardization yields $A_{\text{high}}=+1$ and $A_{\text{low}}=-1$; if the rewards are equal we set $A_{1}=A_{2}=0$. Thus, for any ordered pair $(y^{(1)},y^{(2)})$ with distinct rewards,
$g(\theta)=\tfrac{1}{GH}\left(S(y^{(1)})-S(y^{(2)}))\cdot\mathrm{sign}(r(y^{(1)})-r(y^{(2)})\right).$
GRPO.
In GRPO, $r(2,2)=1$ and $r(1,1)=r(2,1)=0$. Thus, the only nonzero contributions come from mixed pairs $\{(2,2),(1,1)\}$ and $\{(2,2),(2,1)\}$. Using independence of the two samples and summing both orderings,
$\bar{g}_{\mathrm{GRPO}}(\theta)=\mathbb{E}[g(\theta)]=\tfrac{2}{GH}\left(\mathbb{P}(2,2)\mathbb{P}(1,1)(S(2,2)-S(1,1))+\mathbb{P}(2,2)\mathbb{P}(2,1)(S(2,2)-S(2,1))\right).$
With $G=H=2$, $\frac{2}{GH}=\frac{1}{2}$. Substituting the probabilities and $S(\cdot)$ yields
$\bar{g}_{\mathrm{GRPO}}(\theta)=\tfrac{1}{2}\begin{bmatrix}p(p-1)q\\ p(1-p)q\\ pq(q-1)\\ pq(1-q)\end{bmatrix}$.
SGPO.
In SGPO, $r_{\mathrm{SGPO}}(2,2)=1$, $r_{\mathrm{SGPO}}(2,1)=\tfrac{1}{2}$, and $r_{\mathrm{SGPO}}(1,1)=0$. Therefore, all three unordered mixed pairs contribute: $\{(2,2),(2,1)\}$, $\{(2,2),(1,1)\}$, and $\{(2,1),(1,1)\}$. Summing both orderings, we have
$\bar{g}_{\mathrm{SGPO}}(\theta)=\tfrac{2}{GH}\left(\mathbb{P}(2,2)\mathbb{P}(2,1)(S(2,2)-S(2,1))+\mathbb{P}(2,2)\mathbb{P}(1,1)(S(2,2)-S(1,1))+\mathbb{P}(2,1)\mathbb{P}(1,1)(S(2,1)-S(1,1)))\right).$
Again, we have $\frac{2}{GH}=\frac{1}{2}$ for $G=H=2$, and substituting gives
$\bar{g}_{\mathrm{SGPO}}(\theta)=\tfrac{1}{2}\begin{bmatrix}p(p-1)\\ p(1-p)\\ p^{2}q(q-1)\\ p^{2}q(1-q)\end{bmatrix}$.
Now, we derive the update rules in Eq. 3 and the functions in Eq. 4 from the population gradients. Let $\theta^{(k+1)}=\theta^{(k)}+\eta\,\bar{g}(\theta^{(k)})$, where $\bar{g}$ is either $\bar{g}_{\mathrm{SGPO}}$ or $\bar{g}_{\mathrm{GRPO}}$. For notational brevity, define the step-1 logits $(\theta_{1},\theta_{2})=(\theta_{x,1}^{1},\theta_{x,2}^{1})$ and step-2 logits $(\theta_{3},\theta_{4})=(\theta_{x,2,1}^{2},\theta_{x,2,2}^{2})$. In addition, we have
$p=\tfrac{e^{\theta_{2}}}{e^{\theta_{1}}+e^{\theta_{2}}},\quad q=\tfrac{e^{\theta_{4}}}{e^{\theta_{3}}+e^{\theta_{4}}}.$
Update of $p$.
Let $g_{1},g_{2}$ denote the first two coordinates of $\bar{g}(\theta)$. Then, we have
$p^{(k+1)}=\tfrac{e^{\theta_{2}^{(k)}+\eta g_{2}}}{e^{\theta_{1}^{(k)}+\eta g_{1}}+e^{\theta_{2}^{(k)}+\eta g_{2}}}=\tfrac{e^{\theta_{2}^{(k)}}e^{\eta g_{2}}}{e^{\theta_{1}^{(k)}}e^{\eta g_{1}}+e^{\theta_{2}^{(k)}}e^{\eta g_{2}}}.$
Using $p^{(k)}=\frac{e^{\theta_{2}^{(k)}}}{e^{\theta_{1}^{(k)}}+e^{\theta_{2}^{(k)}}}$ and $1-p^{(k)}=\frac{e^{\theta_{1}^{(k)}}}{e^{\theta_{1}^{(k)}}+e^{\theta_{2}^{(k)}}}$, we can factor out $e^{\theta_{1}^{(k)}}+e^{\theta_{2}^{(k)}}$ to obtain
$p^{(k+1)}=\tfrac{p^{(k)}e^{\eta(g_{2}-g_{1})}}{(1-p^{(k)})+p^{(k)}e^{\eta(g_{2}-g_{1})}}.$
For SGPO, from $\bar{g}_{\mathrm{SGPO}}(\theta)$ we have $g_{1}=\tfrac{1}{2}p(p-1)$ and $g_{2}=\tfrac{1}{2}p(1-p)$, hence $g_{2}-g_{1}=p(1-p)$. For GRPO, from $\bar{g}_{\mathrm{GRPO}}(\theta)$ we have $g_{1}=\tfrac{1}{2}p(p-1)q$ and $g_{2}=\tfrac{1}{2}p(1-p)q$, hence $g_{2}-g_{1}=p(1-p)q$. Setting $\eta=1$ gives the claimed update
$p^{(k+1)}_{\mathrm{SGPO}}$ $=\tfrac{p^{(k)}_{\mathrm{SGPO}}e^{p^{(k)}_{\mathrm{SGPO}}(1-p^{(k)}_{\mathrm{SGPO}})}}{1-p^{(k)}_{\mathrm{SGPO}}+p^{(k)}_{\mathrm{SGPO}}e^{p^{(k)}_{\mathrm{SGPO}}(1-p^{(k)}_{\mathrm{SGPO}})}}=\exp(f_{11}(p^{(k)}_{\mathrm{SGPO}}))$,
$p^{(k+1)}_{\mathrm{GRPO}}$ $=\tfrac{p^{(k)}_{\mathrm{GRPO}}e^{p^{(k)}_{\mathrm{GRPO}}(1-p^{(k)}_{\mathrm{GRPO}})q^{(k)}_{\mathrm{GRPO}}}}{1-p^{(k)}_{\mathrm{GRPO}}+p^{(k)}_{\mathrm{GRPO}}e^{p^{(k)}_{\mathrm{GRPO}}(1-p^{(k)}_{\mathrm{GRPO}})q^{(k)}_{\mathrm{GRPO}}}}=\exp(f_{21}(p^{(k)}_{\mathrm{GRPO}},q^{(k)}_{\mathrm{GRPO}}))$.
Taking $\log$ on both sides yields
$f_{11}(p)$ $=\log p+p(1-p)-\log\left(1-p+pe^{p(1-p)}\right),$
$f_{21}(p,q)$ $=\log p+p(1-p)q-\log\left(1-p+pe^{p(1-p)q}\right).$
Update of $q$.
An identical argument applies to $q$ using the last two coordinates $g_{3},g_{4}$ of $\bar{g}(\theta)$:
$q^{(k+1)}=\tfrac{q^{(k)}e^{\eta(g_{4}-g_{3})}}{(1-q^{(k)})+q^{(k)}e^{\eta(g_{4}-g_{3})}}.$
For SGPO, $g_{3}=\tfrac{1}{2}p^{2}q(q-1)$ and $g_{4}=\tfrac{1}{2}p^{2}q(1-q)$, so $g_{4}-g_{3}=p^{2}q(1-q)$. For GRPO, $g_{3}=\tfrac{1}{2}pq(q-1)$ and $g_{4}=\tfrac{1}{2}pq(1-q)$, so $g_{4}-g_{3}=pq(1-q)$. With $\eta=1$ this gives
$q^{(k+1)}_{\mathrm{SGPO}}=\exp(f_{12}(p^{(k)}_{\mathrm{SGPO}},q^{(k)}_{\mathrm{SGPO}})),\quad q^{(k+1)}_{\mathrm{GRPO}}=\exp(f_{22}(p^{(k)}_{\mathrm{GRPO}},q^{(k)}_{\mathrm{GRPO}}))$,
where by taking $\log$ on both sides, we have
$f_{12}(p,q)$ $=\log q+p^{2}q(1-q)-\log\left(1-q+qe^{p^{2}q(1-q)}\right),$
$f_{22}(p,q)$ $=\log q+pq(1-q)-\log\left(1-q+qe^{pq(1-q)}\right).$
@@ Appendix B Missing Proofs
We then provide several technical lemmas that are important to the proof of Theorem 3.3. Indeed, the first lemma summarizes the properties of particular functions related to the aforementioned functions $f_{11}$, $f_{21}$, $f_{12}$ and $f_{22}$ from Eq. 4.
Lemma B.1.
@@
    The function $\varphi(x):=\log(1+e^{-e^{x}})$ is strictly concave on $(-\infty,0)$.
@@
$f^{\prime}_{11}(p)=\tfrac{1+(1-2p)p(1-p)}{p(1-p+pe^{p(1-p)})}>\tfrac{3}{4p(1-p+pe^{p(1-p)})}>0.$
@@
$\tfrac{\partial f_{21}(p,q)}{\partial p}$ $=$ $\tfrac{1+q(1-2p)p(1-p)}{p(1-p+pe^{qp(1-p)})}\ >\ \tfrac{3}{4p(1-p+pe^{p(1-p)})}>0,$
$\tfrac{\partial f_{21}(p,q)}{\partial q}$ $=$ $\tfrac{p(1-p)^{2}}{1-p+pe^{qp(1-p)}}\ >\ 0.$
@@
$\varphi^{\prime\prime}(x)=\tfrac{(e^{x+e^{x}}-e^{e^{x}}-1)e^{x}}{e^{2e^{x}}+2e^{e^{x}}+1}.$
@@
$(ue^{u}-e^{u}-1)u=(e^{u}(u-1)-1)u<-u<0.$
Thus, $\varphi^{\prime\prime}(x)<0$ for all $x<0$ which shows that $\varphi$ is strictly concave on $(-\infty,0)$. $\Box$
@@
By Lemma B.3, this is indeed true. This completes the proof. $\Box$
Lemma B.3.
@@
where $c>0$ is a universal constant, and single-variable polynomials $c_{20}(s),\ldots,c_{2}(s),c_{0}(s)>0$ and $\Delta_{2}:=c_{1}(s)^{2}-4c_{2}(s)c_{0}(s)<0$, for all $s>0$ (see Appendix B for details). Notice that from the table, we can see the only nontrivial parts are $c_{3}(s)>0$ and $c_{2}(s)>0$ because only these two contain negative coefficients. The positivity of $c_{3}(s)$ is simple because there is only one term ($s^{9}$) with negative coefficient and for all $s>0$,
@@
Finally, we have $p_{\texttt{SGPO}}^{(0)}\geq q_{\texttt{SGPO}}^{(0)}$. Thus, it suffices to show that $p_{\texttt{SGPO}}^{(k)}\geq q_{\texttt{SGPO}}^{(k)}$ implies $p_{\texttt{SGPO}}^{(k+1)}>q_{\texttt{SGPO}}^{(k+1)}$ for all $k\geq 0$. Indeed, Lemma B.1(i) and $p_{\texttt{SGPO}}^{(k)}\geq q_{\texttt{SGPO}}^{(k)}$ yield
@@
Then, Lemma B.1(ii) and $p_{\texttt{SGPO}}^{(k)}\in(0,1)$ yield